OpenStudy (anonymous):

on a blueprint a line 4 inches long represents 32ft. the number of inches represents 80ft is: a.6.5 b. 8 c.12 d.16 e.none of the above

OpenStudy (anonymous):

e

OpenStudy (anonymous):

how did u try 2 solve it?

OpenStudy (anonymous):

i have a brain u know

OpenStudy (anonymous):

ya i know but i want to know b/c i tried 2 solve itt n got like 3 diff answers

OpenStudy (anonymous):

ya he is correct

OpenStudy (anonymous):

man its called guessing

OpenStudy (anonymous):

but how cuz i gotta teach my brother

OpenStudy (anonymous):

look man just guess how hard can that be

OpenStudy (anonymous):

32/4=8 80/x=8 80/8=x x=10

OpenStudy (anonymous):

kk thx

OpenStudy (anonymous):

i guess it can cost u alot like a D

OpenStudy (anonymous):

dont listen to rld613 he gives wrong answers trust me i know he gave me wrong and i faild

OpenStudy (anonymous):

i think i have rarely answered ur questions

OpenStudy (anonymous):

actually someone else also gave me the same answer

OpenStudy (anonymous):

poor yuna :(

OpenStudy (anonymous):

i gave u the method to how i figured it out and if u want u can follow it otherwise dump it

OpenStudy (anonymous):

Pelaski unfortunately i had to report u since this is an educational site and i wldnt want my kids to be reading this

OpenStudy (anonymous):

pelaski1:lolz thx 4 the warning n dont worry i always study it first b4 believing it rld613: thx i will..u have kids? oh~~ cool .. they are lucky then, u can help w/ their hw

OpenStudy (anonymous):

yeah dude study just warning u about diz dude told u man now he wants to suspend me lol.. u see

OpenStudy (anonymous):

I just want u to watch ur language there are many kids on here and it isnt correct

OpenStudy (anonymous):

i dint even say nothing meat ball head

OpenStudy (anonymous):

if u need a private tutor go ahead and dish out the money. I see you are pained by the mark you recieved and you think it is imp to you so maybe ur parents will understand your concerns

OpenStudy (anonymous):

um... thx both of u but this is getting too..intense...i'm out but a suggestion, lets not get too hype

OpenStudy (anonymous):

see man i was right about him lol.

OpenStudy (anonymous):

u really are wrong abt me!!!!!!

OpenStudy (anonymous):

i wasant talking about u but alright lmao

OpenStudy (anonymous):

I am just offering suggestions my parents wld offer me tutors when i got too stressed out and that is the best option but whatever it is ur perogative

OpenStudy (anonymous):

We r all humans and i give out the wrong answrs too but some of them here like james satellite and myinaya are really good like they teach math